# Why Low-Precision Transformer Training Fails: An Analysis on Flash Attention

### v1

@@ Abstract
The pursuit of computational efficiency has driven the adoption of low-precision formats for training transformer models. However, this progress is often hindered by notorious training instabilities. This paper provides the first mechanistic explanation for a long-standing and unresolved failure case where training with flash attention in low-precision settings leads to catastrophic loss explosions. Our in-depth analysis reveals that the failure is not a random artifact but caused by two intertwined phenomena: the emergence of similar low-rank representations within the attention mechanism and the compounding effect of biased rounding errors inherent in low-precision arithmetic. We demonstrate how these factors create a vicious cycle of error accumulation that corrupts weight updates, ultimately derailing the training dynamics. To validate our findings, we introduce a minimal modification to the flash attention that mitigates the bias in rounding errors. This simple change stabilizes the training process, confirming our analysis and offering a practical solution to this persistent problem. Code is available at https://github.com/ucker/why-low-precision-training-fails.
@@ 3.3 The Root Causes of Training Failure
In these cases, a green $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathtt{1}$} indicates that rounding up is required, whereas a $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathtt{0}$} indicates no rounding is needed. Because the operands $\bar{{\mathbf{P}}}[T,t]{\mathbf{V}}[t,i]$ are negative and have a large exponent (because $\bar{{\mathbf{P}}}[T,t]=1$ does not make $\bar{{\mathbf{P}}}[T,t]{\mathbf{V}}[t,i]$ smaller), even a small rounding error is magnified, resulting in a significant negative error in the final sum. This systematic negative bias in the computation of $\bar{{\mathbf{O}}}$ is the ultimate source of the training failure.
@@
We consider the addition of two negative values, $\bar{{\mathbf{P}}}[T,t_{1}]{\mathbf{V}}[t_{1},i]$ and $\bar{{\mathbf{P}}}[T,t_{2}]{\mathbf{V}}[t_{2},i]$, from Fig. 6. Their BF16 values are computed in FP32 and then rounded. For this example, we start with their exact FP32 representations, which have zero bits beyond the 7th fraction bit:
$\mathtt{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000000}}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0011010}}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}}$ $\;(-2.40625)$ $\mathtt{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000000}}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010011}}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}}$ $\;(-2.296875)$
@@
$(\mathtt{-1.}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0011010}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}})+(\mathtt{-1.}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010011}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}})$ $=$ $-\mathtt{10.{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0101101}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}}$
@@
The exact result in FP32 is $\mathtt{1}\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000001}}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010110}}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathtt{1}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{000000000000000}}$, which corresponds to 4.703125-4.703125. To store this result in BF16, it must be rounded to 7 fraction bits. The significand is $\mathtt{1.{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0010110}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}1}}$. The 7-bit fraction is $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010110}$}, and the first truncated bit is $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathtt{1}$}. According to the “round to nearest, ties to even” rule, this requires rounding up (adding 1 to the last bit of the fraction):
@@
The final BF16 result is $\mathtt{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000001}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010111}}$, which represents 4.71875-4.71875. This rounded value is more negative than the true sum of 4.703125-4.703125. The error introduced by this single addition is $\mathbf{-0.015625}$. When such rounding events occur systematically across many additions in the $\bar{{\mathbf{P}}}{\mathbf{V}}$ product, the errors accumulate, creating the negative bias in $\mathbf{O}$} that ultimately destabilizes training.
@@
With more than one $\bar{{\mathbf{P}}}[T,t]=1$ and negative ${\mathbf{V}}[t,i]$, the addition in $\bar{{\mathbf{P}}}{\mathbf{V}}$ can cause significand overflow. This necessitates a right shift, which often forces a round-down, introducing negative rounding error in $\mathbf{O}$} and leading to positive $({\bm{\delta}}_{lp}-{\bm{\delta}}_{hp})[T]$.

### v2

@@ Abstract
The pursuit of computational efficiency has driven the adoption of low-precision formats for training transformer models. However, this progress is often hindered by notorious training instabilities. This paper provides the first mechanistic explanation for a long-standing and unresolved failure case where training with flash attention in low-precision settings leads to catastrophic loss explosion. Our in-depth analysis reveals that the failure is not a random artifact but caused by two intertwined phenomena: the emergence of similar low-rank representations within the attention mechanism and the compounding effect of biased rounding errors inherent in low-precision arithmetic. We demonstrate how these factors create a vicious cycle of error accumulation that corrupts weight updates, ultimately derailing the training dynamics. To validate our findings, we introduce a minimal modification to the flash attention that mitigates the bias in rounding errors. This simple change stabilizes the training process, confirming our analysis and offering a practical solution to this persistent problem. Code is available at https://github.com/ucker/why-low-precision-training-fails.
@@ 3.3 The Root Causes of Training Failure
Because the summation is performed in FP32, the accumulation of lower-order bits from the small numbers can activate the sticky bit. Consequently, this forces a rounding up when the subsequent BF16 number is added. Therefore, the rounding bit $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathtt{1}$} indicates that rounding up is required. Because the operands $\bar{{\mathbf{P}}}[T,t]{\mathbf{V}}[t,i]$ are negative and have a large exponent (because $\bar{{\mathbf{P}}}[T,t]=1$ does not make $\bar{{\mathbf{P}}}[T,t]{\mathbf{V}}[t,i]$ smaller), the error of rounding up is magnified, resulting a negative error. When the rounding bit is $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathtt{0}$}, the result is rounded down, introducing a positive error. Furthermore, the positive error is smaller compared to the negative error from rounding up, as the values being added are usually very small. This asymmetry leads to the rounding error dominated by the rounding up, resulting in negative rounding error as observed in our analysis. This systematic negative bias in the computation of $\bar{{\mathbf{O}}}$ is the ultimate source of the training failure.
@@
To make this concrete, we now analyze the specific BF16 number addition that causes the large negative error jump shown in Fig. 6(c). Because the summation is performed in FP32, the first BF16 value is added by some small values from other tokens before the second BF16 value is added. This can activate the sticky bit, which can force a round-up when adding the second BF16 value. For this example, we start with the FP32 representations. The first operand is the accumulated sum of previous terms, which includes a BF16 ($\mathtt{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000000}}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0011010}}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}};-2.40625$) plus small residual ($\sim 0.00087$) that activate the sticky bit. The second operand is another BF16 value. Their FP32 representations are:
$\mathtt{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000000}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0011010}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000111000101110}}$ $\;(-2.4071154594421387)$ $\mathtt{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000000}}}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010011}}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}}$ $\;(-2.296875)$
@@
$(\mathtt{-1.}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0011010}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000111000101110}})+(\mathtt{-1.}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010011}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000000000000000}})$ $=$ $-\mathtt{10.{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0101101}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\mathtt{0000111000101110}}$
@@
[…]
@@
The final BF16 result is $\mathtt{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathtt{10000001}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathtt{0010111}}$, which represents 4.71875-4.71875. This rounded value is more negative than the true sum of 4.703990459442139-4.703990459442139. The error introduced by this single addition is $\mathbf{-0.014759540557861328}$. When such rounding events occur systematically across many additions in the $\bar{{\mathbf{P}}}{\mathbf{V}}$ product, the errors accumulate, creating the negative bias in $\mathbf{O}$} that ultimately destabilizes training.
@@
With more than one $\bar{{\mathbf{P}}}[T,t]=1$ and negative ${\mathbf{V}}[t,i]$, the addition in $\bar{{\mathbf{P}}}{\mathbf{V}}$ can cause significand overflow. This necessitates a right shift and a significand round-up due to sticky bit, introducing negative rounding error in $\mathbf{O}$} and leading to positive $({\bm{\delta}}_{lp}-{\bm{\delta}}_{hp})[T]$.